OpenStudy (anonymous):

how to find the integral of 1/((1/x)+tan(x))

OpenStudy (anonymous):

I only got as far as int (xcosx)/(cosx +xsinx). From there I cannot figure a sub that would work . let the denominator = (1/x + sinx/cosx) = cosx +xsinx / xcosx which inverses to what I got. From there, it's got me.

OpenStudy (anonymous):

maybe you can figure someway, but that's got it moving a little..

OpenStudy (anonymous):

Try u sub; u = 1/x + tan x